In a carnival game a single die is rolled. If the number that shows on top is odd, you win that amount. If the number showing on top is even, you lose that amount. Construct a probability distribution for this game and find the expected value for a person playing this game.
I have done a few of these, but this one throws me.

Answer by robertb(5830)

Let X be the random variable representing the earnings/loss at a single throw. Then X = 1,3,5,-2,-4,-6. We get the following discrete uniform probability mass function:
P(1) = P(2) =P(3) = P(-2) = P(-4) = P(-6) = 1/6.
The expected value is then .